# The Evolution of Reinforcement Learning in Quantitative Finance

### v1

@@ 5.1. Introduction
In this section, we emphasise two pivotal facets of the RL framework: the action space and the reward functions, which are both contingent on the specific application. Actions, denoted as $A_{t}$, are the input of the agent into the environment at time $t$. Typically, in the QF-related RL literature, agents are assumed to have no impact on the environment, and actions may be discrete, such as buying or selling a stock, or continuous, such as portfolio weights at time $t$. Table 3 summarises the action and reward functions used in the fifteen most cited papers in this survey, which are also chronologically ordered and will be discussed further in this section.
@@
Drawing parallels with finance, this approach conceptually resembles the discounting of the price of bonds, where $\gamma$ is the period discount factor and $R_{i}$ represents the coupon plus the face value at time $i$ (FABOZZI_2012, ). In finite-time horizons, such as modelling the expected return until a terminal event such as bankruptcy, using a discount factor $\gamma<1$ is consistent with financial models such as the Gordon Growth Model (GORDON_1959, ) and bond pricing. These models inherently apply a discount factor of less than 1 to reflect the time value of money [^18] There are cases where the time value of money principle is violated, such as with an inverted yield curve, where short-term rates exceed long-term rates, making future cash flows more valuable. In RL, this could correspond to $\gamma>1$, where future rewards are prioritised. This is uncommon in RL because it can lead to instability in learning algorithms and unbounded returns, complicating optimisation., ensuring that the present value of future cash flows is properly calculated. This method emphasises the financial principle that future rewards are typically less valuable than immediate rewards, even within a finite horizon. However, setting $\gamma=1$ can be appropriate and meaningful in specific bounded scenarios where future rewards are equally critical.
Neuneier (NEUNEIER_1996, ) FX/Equities 2 x x
Dempster et al. (DEMPSTER_2001, ) FX 2, 3 x x
Nevmyvaka et al. (NEVMYVAKA_2006, ) Equities 11 x x
Dempster and Leemans (DEMPSTER_2006, ) FX 2 Tanh x x
Lee et al. (LEE_2007, ) Equities 2-7 x x
Deng et al. (DENG_2016, ) Com./ Eq. Index 3 Tanh x x
Jiang and Liang (JIANG_2017, ; JIANG_2017_1, ) Crypto 12 x Softmax x x
Almahdi and Yang (ALMAHDI_2017, ) Stock ETF’s 5 x Logsig/Softmax x x
Liang et al. (LIANG_2018, ) Equities 5 x x x
Xiong et al. (XIONG_2018, ) Equities 3 x
Jeong and Kim (JEONG_2019, ) Equity Index 3 x
@@ 5.3. Rewards
Most academic research gravitates towards profit-based rewards (CHAN_2001, ; EILERS_2014, ; KAUR_2017, ; SI_2017, ; SORMAYURA_2019, ; SAWHNEY_2021, ). Risk Performance Measures (RPM), such as SR, comprise the second most common focus (BERTOLUZZO_2012, ; CORAZZA_2015, ; WANG_2019, ; THEATE_2021, )). An intriguing variation emerges in several studies (MARINGER_2014, ; ZHANG_2014, ; GABRIELSSON_2015, ; LIU_2020, ; LI_2021, ), notably the use of Differential SR (DSR) as introduced and developed further by Moody and colleagues (MOODY_1997, ; MOODY_1998, ; MOODY_1998_2, ; MOODY_2001, ). DSR, an online variant of standard SR, offers several benefits, including enabling efficient online optimisation and weighing more recent returns (MOODY_1997, ). The generic derivation of the DSR suggests the feasibility of online versions for other performance measures (MOODY_1997, ; BERTOLUZZO_2007, ; ALMAHDI_2017, ; ALMAHDI_2019, ). Alternative measures have been used in various contexts, for instance, Carapuço et al. (CARAPUCO_2018, ) and Wu et al. (WU_2020, ) use the Sortino ratio [^19] The Sortino ratio is similar to the SR but penalises those returns below a given investor threshold (SORTINO_1994, ).. The diversity of these measures highlights the potential to explore broader reward functions, such as incorporating higher moments into the classic SR (PEZIER_2006, ).
@@ 7.1. Introduction
This section reviews common modern finance applications and RL-based solutions and examines RL framework customisation for classic QF issues. The literature illustrates RL’s adaptability to the financial domain as can be seen in the range of popular applications listed in Table 4. In the following, we provide an examination of the salient ideas and potential extensions within the current literature on this topic and provide future directions in these areas. In addition, we delineate the pivotal concepts and thematic trends that have emerged. Although we acknowledge the extensive literature on PM, we have reserved some works for inclusion in different subsections as they emphasise distinct methods or themes relevant to PM. For example, the papers by Ding et al. (DING_2018, ) and Yu et al. (YU_2019, ), which use IL methods within the PM environment, are already discussed in subsection 6.5.
Publication Applications
Neuneier (NEUNEIER_1996, ) x x
Moody and Wu (MOODY_1997, ) x
Moody et al. (MOODY_1998, )
Moody and Safell (MOODY_2001, )
Dempster et al. (DEMPSTER_2001, ) x
Nevmyvaka et al. (NEVMYVAKA_2006, ) x
Dempster and Leemans (DEMPSTER_2006, ) x x
Lee et al. (LEE_2007, ) x x
Deng et al. (DENG_2016, ) x
Jiang and Liang (JIANG_2017, ) x x
Jiang and Liang (JIANG_2017_1, )
Almahdi and Yang (ALMAHDI_2017, ) x x
Liang et al. (LIANG_2018, ) x x
Xiong et al. (XIONG_2018, ) x
Jeong and Kim (JEONG_2019, ) x
@@ 7.7. Business Cycles and Augmenting Existing Trading Strategies
Kim and Kim (KIM_2019, ) implemented a DQN for pair trading, rewarding the agent when the spread between two similar companies reaches a predefined threshold and reverts to the mean. The agent is penalised if stop-loss limits are hit or the spread fails to revert. The environment focuses solely on the spread but could be enhanced by incorporating factors like liquidity. Wang et al. (WANG_C_2021, ) applied a similar strategy to five-year data from 75 NASDAQ companies, outperforming the baseline in terms of SR. Cartea et al. (CARTEA_2021, ) are perhaps the first to apply DDQN and Reinforced Deep Markov Models (RDMMs) (FERREIRA_2020, ) to FX triplets. In an FX triplet, one of the three FX pairs is redundant due to the no-arbitrage rule in frictionless markets [^24] The no-arbitrage rule asserts that in an efficient market, it is impossible to make a risk-free profit through simultaneous buying and selling of related assets, as price discrepancies are quickly corrected.. Therefore, establishing an opposing position is expected to generate profit when the triplet deviates from the no-arbitrage relationship. Through a simulation study, the authors favoured the use of RDMMs. Kim et al. (KIM_2022, ) proposed a hybrid RL framework in which the first component, based on the TD3 algorithm (Actor-Critic) (FUJIMOTO_2018, ), is responsible for trading actions in pairs trading, while the second component, based on a DDQN agent, is responsible for stop-loss boundaries. In recognition of the risk of a structural break in the trading pair, Lu et al. (LU_2022, ) introduced a hybrid RL-based on DQN in which structural break detection mechanisms are used as input to a DQN agent.
@@ 7.9. Human in the Loop – Robo-Advising
Most applications reviewed here aim to minimise human intervention and implement autonomous trading systems. However, Robo-advising stands out as an area where human input remains essential. Characterised as a class of algorithms, Robo-advisors use client-specific risk appetites for financial planning and investment advice, offering an innovative alternative to traditional investment management. Human input is necessary for establishing risk preferences. Robo-advisors confer multiple advantages over traditional investment managers, including cost-effectiveness due to reduced overheads, such as office maintenance. In addition, they mitigate human biases associated with asset allocation, recommendations, and data collection (FOERSTER_2017_A, ). Moreover, Robo-advising has found acceptance among systemically significant entities, indicating its growing industry significance. Historically, risk preferences were determined through questionnaires (CHANRNESS_2013, ), a method potentially subject to bias (HOLT_2002, ). In contrast, the studies in this section use IRL and Inverse Portfolio Optimisation (IPO) to estimate client risk tolerance.
Central to the reviewed literature is the concept that Robo-advisors learn or infer an investor’s risk preferences through iterative interactions. For instance, in the work of Alsabah et al. (ALSABAH_2021, ), the Robo-advisor allocates the investor’s capital across several pre-constructed portfolios, each reflecting distinct levels of risk preference. These preferences can then be updated when the Robo-advisor requests the investor to modify portfolio allocations. This methodology, while innovative, incurs opportunity costs, as investors must spend time understanding the financial environment in which their wealth is invested. Furthermore, this framework introduces an exploration/exploitation trade-off. Robo-advisors must choose between investing based on current risk preference estimates or seeking updated risk preferences, which may be costly. The variability of investor preferences over time adds another layer of complexity. To address these challenges, the authors propose a planning algorithm similar to the IRL framework and demonstrate that convergence to the optimal policy correlates with the consistency of investor choices. Further innovation comes from Wang and Yu (WANG_2021_D, ), who proposed a two-agent framework: the first agent uses IPO to infer investor preferences and expected returns from past allocations, while the second agent, based on DDPG, performs a multi-period mean-variance optimisation. Yu and Dong (YU_2020, ) also use IPO, with the DDPG agent serving as a benchmark case, while Dixon and Halperin (DIXON_2020, ) apply G-Learning (FOX_2015, ) in an IRL context, with applications in wealth management and Robo-advising.
Despite the intrigue surrounding Robo-advising, the field remains under-researched. As one of the few RL applications in finance involving human feedback, future research could explore multi-agent Robo-advisors that compete or collaborate. This could illuminate how risk preferences operate in this context and reveal aggregate risk preferences.
@@ 8. Discussion
Another issue to address is the evident survivorship bias [^26] Refers to the selection of stocks that have performed well and survived throughout the sample period. in some publications, for example, Kaur (KAUR_2017, ) and Wu et al. (WU_2020, ). These works often apply their frameworks to successful single stock names like Apple Inc., neglecting to test their methodology on stocks that are currently non-existent or bankrupt. Furthermore, the focus on a limited set of tradable assets introduces potential survivorship bias, and the success of these assets does not necessarily imply effectiveness across other available assets. To lend greater authenticity, we should scrutinise the proposed trading system on benchmark constituents over time. Recent studies have shown progress in this regard. For example, Lee et al. (LEE_2020, ) extended their RL framework to the Russell 3000, while Wang et al. (WANG_2019, ) used Wharton Research Data Services (WRDS) to select a quarter of valid stocks annually.
@@ 9. Future Directions
This section outlines promising research opportunities emerging from our comprehensive analysis of RL in QF:

### v2

@@ 5.1. Introduction
In this section, we emphasise two pivotal facets of the RL framework: the action space and reward functions tailored to finance. Actions, denoted $A_{t}$, represent the agent’s interactions with the environment at time $t$. In QF-related RL studies, agents typically do not affect the environment, with actions ranging from discrete (e.g., buy or sell decisions) to continuous (e.g., adjusting portfolio weights). Table 3 (in Appendix) presents the action and reward functions from the top fifteen cited papers in our survey.
@@
Drawing parallels with finance, this approach conceptually resembles the discounting of the price of bonds, where $\gamma$ is the period discount factor and $R_{i}$ represents the coupon plus the face value at time $i$ (FABOZZI_2012, ). In finite-time horizons, such as modelling the expected return until a terminal event such as bankruptcy, using a discount factor $\gamma<1$ is consistent with financial models such as the Gordon Growth Model (GORDON_1959, ) and bond pricing. These models inherently apply a discount factor of less than 1 to reflect the time value of money, ensuring that the present value of future cash flows is properly calculated. This method emphasises the financial principle that future rewards are typically less valuable than immediate rewards, even within a finite horizon. However, setting $\gamma=1$ can be appropriate and meaningful in specific bounded scenarios where future rewards are equally critical.
@@ 5.3. Rewards
Most academic research gravitates towards profit-based rewards (CHAN_2001, ; EILERS_2014, ; KAUR_2017, ; SI_2017, ; SORMAYURA_2019, ; SAWHNEY_2021, ). Risk Performance Measures (RPM), such as SR, comprise the second most common focus (BERTOLUZZO_2012, ; CORAZZA_2015, ; WANG_2019, ; THEATE_2021, )). An intriguing variation emerges in several studies (MARINGER_2014, ; ZHANG_2014, ; GABRIELSSON_2015, ; LIU_2020, ; LI_2021, ), notably the use of Differential SR (DSR) as introduced and developed further by Moody and colleagues (MOODY_1997, ; MOODY_1998, ; MOODY_1998_2, ; MOODY_2001, ). DSR, an online variant of standard SR, offers several benefits, including enabling efficient online optimisation and weighing more recent returns (MOODY_1997, ). The generic derivation of the DSR suggests the feasibility of online versions for other performance measures (MOODY_1997, ; BERTOLUZZO_2007, ; ALMAHDI_2017, ; ALMAHDI_2019, ). Alternative measures have been used in various contexts, for instance, Carapuço et al. (CARAPUCO_2018, ) and Wu et al. (WU_2020, ) use the Sortino ratio (SORTINO_1994, ). The diversity of these measures highlights the potential to explore broader reward functions, such as incorporating higher moments into the classic SR (PEZIER_2006, ).
@@ 7.1. Introduction
This section reviews common modern finance applications and RL-based solutions and examines RL framework customisation for classic QF issues. The literature illustrates RL’s adaptability to the financial domain as can be seen in the range of popular applications listed in Table 4 (in Appendix). In the following, we provide an examination of the salient ideas and potential extensions within the current literature on this topic and provide future directions in these areas. In addition, we delineate the pivotal concepts and thematic trends that have emerged. Although we acknowledge the extensive literature on PM, we have reserved some works for inclusion in different subsections as they emphasise distinct methods or themes relevant to PM. For example, the papers by Ding et al. (DING_2018, ) and Yu et al. (YU_2019, ), which use IL methods within the PM environment, are already discussed in subsection 6.5.
@@ 7.7. Business Cycles and Augmenting Existing Trading Strategies
Kim and Kim (KIM_2019, ) implemented a DQN for pair trading, rewarding the agent when the spread between two similar companies reaches a predefined threshold and reverts to the mean. The agent is penalised if stop-loss limits are hit or the spread fails to revert. The environment focuses solely on the spread but could be enhanced by incorporating factors like liquidity. Wang et al. (WANG_C_2021, ) applied a similar strategy to five-year data from 75 NASDAQ companies, outperforming the baseline in terms of SR. Cartea et al. (CARTEA_2021, ) are perhaps the first to apply DDQN and Reinforced Deep Markov Models (RDMMs) (FERREIRA_2020, ) to FX triplets. In an FX triplet, one of the three FX pairs is redundant due to the no-arbitrage rule in frictionless markets. Therefore, establishing an opposing position is expected to generate profit when the triplet deviates from the no-arbitrage relationship. Through a simulation study, the authors favoured the use of RDMMs. Kim et al. (KIM_2022, ) proposed a hybrid RL framework in which the first component, based on the TD3 algorithm (Actor-Critic) (FUJIMOTO_2018, ), is responsible for trading actions in pairs trading, while the second component, based on a DDQN agent, is responsible for stop-loss boundaries. In recognition of the risk of a structural break in the trading pair, Lu et al. (LU_2022, ) introduced a hybrid RL-based on DQN in which structural break detection mechanisms are used as input to a DQN agent.
@@ 7.9. Human in the Loop – Robo-Advising
Most applications discussed aim to reduce human intervention by implementing autonomous trading systems, with Robo-advising as a notable exception where human input is crucial. Defined as algorithms that cater to client-specific risk appetites, Robo-advisors offer a cost-effective alternative to traditional investment management, minimising overheads and human biases in asset allocation and recommendations (FOERSTER_2017_A, ). Historically, risk preferences were gauged through potentially biased questionnaires (CHANRNESS_2013, ; HOLT_2002, ), whereas contemporary approaches like Inverse Reinforcement Learning (IRL) and Inverse Portfolio Optimisation (IPO) dynamically estimate these preferences (ALSABAH_2021, ; WANG_2021_D, ; YU_2020, ).
The reviewed studies highlight that Robo-advisors iteratively learn an investor’s risk tolerance by adjusting portfolio allocations based on the investor’s feedback. This interactive process, however, introduces a trade-off between exploiting known preferences and exploring potentially more optimal allocations, adding complexity due to the variability of investor preferences over time (ALSABAH_2021, ). Innovations include a dual-agent framework where one agent infers preferences and expected returns, while the other optimises investments over multiple periods (WANG_2021_D, ). Further applications in wealth management use advanced algorithms like G-Learning to refine these methods (DIXON_2020, ). Despite Robo-advising’s growing industry significance, it remains under-researched. Future studies could focus on multi-agent systems in Robo-advising to understand competitive or cooperative dynamics in risk preference management, potentially revealing broader implications for investor behaviour.
@@ 8. Discussion
Another issue to address is the evident survivorship bias in some publications, for example, Kaur (KAUR_2017, ) and Wu et al. (WU_2020, ). These works often apply their frameworks to successful single stock names like Apple Inc., neglecting to test their methodology on stocks that are currently non-existent or bankrupt. Furthermore, the focus on a limited set of tradable assets introduces potential survivorship bias, and the success of these assets does not necessarily imply effectiveness across other available assets. To lend greater authenticity, we should scrutinise the proposed trading system on benchmark constituents over time. Recent studies have shown progress in this regard. For example, Lee et al. (LEE_2020, ) extended their RL framework to the Russell 3000, while Wang et al. (WANG_2019, ) used Wharton Research Data Services (WRDS) to select a quarter of valid stocks annually.
@@ Acknowledgements.
APPENDIX
Publication Data Freq. Period PH TI CP ME PV CH OB Oth. FM
Neuneier (NEUNEIER_1996, ) Daily 1986-1996 x x x x
Moody and Wu (MOODY_1997, ) Monthly/30-min. Several x x
Moody et al. (MOODY_1998, ) Monthly/30-min. Several x x
Moody and Safell (MOODY_2001, ) Monthly/30-min. Several x x
Dempster et al. (DEMPSTER_2001, ) 1-min. 1994-2002 x x x
Dempster and Leemans (DEMPSTER_2006, ) 1-min. 2000-2002 x x x
Lee et al. (LEE_2007, ) Daily 1999-2005 x x x x
Deng et al. (DENG_2016, ) Tick-level Several x x x
Jiang and Liang (JIANG_2017, ) 30-min. 2014-2017 x x x x
Jiang and Liang (JIANG_2017_1, ) 30-min. 2014-2017 x x x x
Almahdi and Yang (ALMAHDI_2017, ) Weekly 2011-2015 x x
Liang et al. (LIANG_2018, ) Daily 2015-2017 x x x
Xiong et al. (XIONG_2018, ) Daily 2009-2018 x x x x
Jeong and Kim (JEONG_2019, ) Daily Several x x
Neuneier (NEUNEIER_1996, ) FX/Equities 2 x x
Dempster et al. (DEMPSTER_2001, ) FX 2, 3 x x
Nevmyvaka et al. (NEVMYVAKA_2006, ) Equities 11 x x
Dempster and Leemans (DEMPSTER_2006, ) FX 2 Tanh x x
Lee et al. (LEE_2007, ) Equities 2-7 x x
Deng et al. (DENG_2016, ) Com./ Eq. Index 3 Tanh x x
Jiang and Liang (JIANG_2017, ; JIANG_2017_1, ) Crypto 12 x Softmax x x
Almahdi and Yang (ALMAHDI_2017, ) Stock ETF’s 5 x Logsig/Softmax x x
Liang et al. (LIANG_2018, ) Equities 5 x x x
Xiong et al. (XIONG_2018, ) Equities 3 x
Jeong and Kim (JEONG_2019, ) Equity Index 3 x
Publication Applications
Neuneier (NEUNEIER_1996, ) x x
Moody and Wu (MOODY_1997, ) x
Moody et al. (MOODY_1998, )
Moody and Safell (MOODY_2001, )
Dempster et al. (DEMPSTER_2001, ) x
Nevmyvaka et al. (NEVMYVAKA_2006, ) x
Dempster and Leemans (DEMPSTER_2006, ) x x
Lee et al. (LEE_2007, ) x x
Deng et al. (DENG_2016, ) x
Jiang and Liang (JIANG_2017, ) x x
Jiang and Liang (JIANG_2017_1, )
Almahdi and Yang (ALMAHDI_2017, ) x x
Liang et al. (LIANG_2018, ) x x
Xiong et al. (XIONG_2018, ) x
Jeong and Kim (JEONG_2019, ) x